The stopping distance d of a car after the brakes are applied varies directly as the square of the speed r. If a car travelling 30 mph can stop in 50 ft, how many feet will it take the same car to stop when it is travelling 20 mph?
Answer by stanbon(75887)

The stopping distance d of a car after the brakes are applied varies directly as the square of the speed r.
---
d = kr^2
---
Solve for "k" using "a car travelling 30 mph can stop in 50 ft"
50 = k*30^2
k = 50/900
k = 0.0556
---------------
Equation:
d = 0.0556*r^2
---
How many feet will it take the same car to stop when it is travelling 20 mph?
d = 0.0556*20^2
---
distance = 22.22 ft
================================
Cheers,
Stan H.
========================